A space vehicle is coasting at a constant velocity of 15.7 m/s in the +y direction relative to a space station. The pilot of the vehicle fires a RCS (reaction control system) thruster, which causes it to accelerate at 0.342 m/s2 in the +x direction. After 34.7 s, the pilot shuts off the RCS thruster. After the RCS thruster is turned off, find (a) the magnitude and (b) the direction of the vehicle's velocity relative to the space station. Express the direction as an angle (in degrees) measured from the +y direction. (Write on paper)
Answer #1

Vx = a*t = 0.342*34.7 = 11.8674 m/s

a) V = √(Vx² + Vy²) = 19.68 m/s relative to station.

b) Θ = arctan[Vx/Vy] = 37.1° right of +y